ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

2. Algebra

1991 MS Classification: 20Mxx, 81Rxx Duplij, Steven, Kharkov State University, Ukraine Continuous bands from supermatrices We investigate continuous supermatrix representations of semigroup bands. We study the Green’s relations on continuous zero semigroups and wreath rectangular bands. We found the “fine” equivalence relations which generalize Green’s relations on subsemigroups in some extent and lead to the “multidimensional” analog of eggbox diagrams. Reference: S. Duplij, Supermatrix representations of semigroup bands, Pure Math. Appl. , 7 (1996) 235–261 (funct-an/9609002).